# Consumption Smoothing in Metropolis: Evidence from the Working-class Households in Prewar Tokyo

### v8

@@ 4.3 Event-Study Analysis: Response to the Advers Health Shocks
Column 1 shows the results of equation 2, in which the head’s earning is replaced with the head’s illness (i.e., the number of days with illness in the previous semi-month cell) as the idiosyncratic shock variable. The family size control, household fixed effects, and time-fixed effects are included in all the regressions. The estimated coefficients on the head’s illness are listed in the subcolumn named “Coef.”.
Column 2 shows the results of equation 2, in which the head’s earning is treated as the endogenous variable. The head’s illness is an instrumental variable (i.e., excluded exogenous regressor) in the first-stage reduced-form equation. The first-stage result is identical to the result listed in Column (2) of Table 6 with the first-stage $F$-statistics of $13.6$. The estimated coefficients on the head’s earning are listed in the subcolumn named “Coef.”.
@@
Panel B: “Other member’s earnings” refers to the total income earned by all family members except for the head of the household. “Sales of assets” includes the sale of daily miscellaneous goods such as newspapers and empty bottles. All dependent variables are in yen.
I test the risk-coping strategies for mitigating adverse health shocks in Table 8. Column 1 lists the reduced-form results for the risk-coping strategies in the same manner in Table 5. The estimated coefficients on the net income variables are positive in most cases, indicating the households’ compensating responses against the shocks. The estimate for the credit purchases category is weakly statistically significant, with the largest magnitude. It suggests that the one-day increase in the head’s illness increased the net credit purchases by $1.29$ yen, which accounts for roughly half of their average daily wages ($1.29/2.56$). Column 2 presents the estimates from the structural equation using the head’s earnings as the endogenous variable. A similar but much clearer result is obtained: the estimate for the credit purchase category is roughly four times greater magnitude than that reported in Column (1) of Table 5. The estimate indicates that a one yen decrease in the head’s earnings increases the net income from credit purchases by $0.55$ yen. This magnitude is consistent with the direct effect of the health shocks on the net credit purchase estimated in Column (1) of Table 8.747474This means that $2.56\times 0.55=1.4$ yen, which is close to $1.3$ yen in Column (1). The estimates for the savings and gifts categories are still negative but now statistically insignificant, supporting the evidence that credit purchase is the primary risk-coping strategy for dealing with health shocks. Since it shall take more time (steps) from withdrawing and receiving gifts to use the money at the retailers, it is likely that the households prefer to use credit purchases.
Overall, the event-study analysis using adverse health shocks confirms baseline findings and provides a sharper illustration of consumption smoothing behaviors in the case of adverse idiosyncratic shock events. In short, my results show that the unique available response among the THBS households to the idiosyncratic income shock was to use credit purchases. The available compensation via the credit purchases accounts for roughly half of their temporary income losses, which partially contributes to smoothing their food consumption in the short-run.
@@ 5 Credit System in Local Economy
Foregoing results shed new light on the roles of credit institutions among skilled workers’ households during historical industrialization. Since the credit purchases are provided by local retailers, this also suggests the importance of the relationships between the workers and retail shops. To the best of my knowledge, however, there is little historical work on credit purchases in prewar Japan. Hence, the details of this credit institution are not yet clear: for example, how retailers get the credit information of their customers, whether the contracts are explicit or implicit, and whether the lenders enforce penalties for the debtors.
@@
Notes: This table summarizes the statistics on the sales in the $326$ retail shops selling white rice, fish, fruit and vegetables, and firewood and charcoal in Tokyo surveyed in 1935. The number of stores reported the share of regular customers (column 2) are $90$, $56$, $61$, and $93$ in the white rice, fish, greengrocer, and fuel (firewood and charcoal) categories, respectively. The “share of sales for regular customers” is defined as the percentage of sales for regular customers in the annual total sales in 1935. The definition of the social classes (“Upper”; “Middle”; “Other”) is not exactly defined in the report. The share of sales method is based on the percentage of sales using each method in the annual total sales in 1935. The number of competitors indicates the number of peers within a three-cho, approximately $327$ meters, radius. The department stores and retail markets are not included in this survey.
@@
The information on the credit purchase is considerably personal, as in the case of state-contingent loans in the rural credit market (Udry 1994), and can hardly be documented. Therefore, systematic information on credit purchases in the prewar period is unavailable. Fortunately, however, a few official survey reports documented some information on this payment method. A survey report on retailers in 1930 (hereafter 1930 Survey Report) revealed that credit purchase was a popular payment method in transactions with consumers. [^75] This survey was conducted in February and March 1930 by Tokyo City and covered $3,892$ small and medium enterprises (Tokyo City Office 1932a, p. 1). Those include not only retailers but also wholesale and other small businesses, and the number of respondents varies depending on the question items. The number of respondents in the statistical table on the payment methods is $2,654$, of which $1,197$ are retailers (Tokyo City Office 1932b, p. 273). To the best of my knowledge, this is the most comprehensive survey of the retailers in Tokyo around 1930. For example, 99%(155/157) of grain retailers, 81%(35/43) of fish retailers, 78%(64/82) of greengrocers, and 94%(45/48) of fuel retailers used credit purchases for their selling method (Tokyo City Office 1932b, pp. 260; 263). A survey report on retail shops in 1935 published by the Tokyo City Chamber of Commerce and Industry provides more detailed information. Table 9 summarizes the information on the sales in those retail shops of rice, fish, fruit and vegetables, and fuels. [^76] This survey covers 939 shops and was conducted on 31st December 1935. I digitized the statistics for all the shops selling the representative items (rice, fish, fruit and vegetables, and fuels) in the credit purchases. The survey instrument includes detailed questions regarding sales and purchases, making it the most comprehensive of the available documents (Tokyo City Chamber of Commerce and Industry 1937, pp. legend, questionnaire). Online Appendix Table C.1 decomposes the statistics listed in Table 9 by the scale of businesses. I confirm that the main findings are materially similar. First, many annual sales are from their regular customers, most of whom are from the non-upper classes. For example, $73$% of annual sales in rice retailers are from their regular customers, and more than $80$% of them are from non-upper classes, including the factory workers. Second, credit purchase is a popular sales method, especially among rice and fuel retailers: more than $70$% of the sales were done with credit in both shops. [^77] These figures are systematically lower than those from the 1930 Survey Report. Given that both surveys show a similar tendency regarding the use of credit purchases, this may be due to the differences in the samples, the definition of the retailers in each category, and the definition of the credit purchase itself. Although cash was also used frequently among greengrocers, $34$% of the sales were settled using credit. This supports the evidence that credit purchase was a popular payment method, and regular customers settled a large part of it. [^78] The 1930 Survey reveals that the retailers had also commonly used credit purchases to buy their source goods from wholesalers (Tokyo City Office 1932a, pp. 154–157). This means that credit purchase was a familiar payment method for retailers as well as consumers. Although these statistics are from the mid-1930s, this confirms the possibility that the retailers were engaged in credit transactions in the context of a long-term relationship with their regular customers. [^79] A textbook for retailers published in 1924 explains an interesting mechanism behind this long-term relationship from an aspect of customer psychology: credit purchase is made not only for the convenience of the customer’s budget but also because of the “vanity” of being able to be a regular customer (Shimizu 1924, pp. 115–116). The final column of this table further provides suggestive evidence behind this. The survey asked retailers whether the sales area was competitive or not, and if yes, they were asked to report the number of competitors within a three-cho (roughly $330$ meters) radius. As shown, rice and fuel stores have 2-3 more competitors than the retailers selling fish, fruit, and vegetables. Given the relatively high utilization of credit in rice and fuel retailers, they may have responded to competition by allowing their customers to trade on credit.
@@
Note: Figure 6a, and 6c illustrate the time-series plots of the average daily and semi-monthly credit purchase and payment, respectively. The blue and red dashed lines indicate the average credit purchases and payments, respectively. Figure 6b and 6d illustrate the time-series plots of the average daily and semi-monthly credit purchase minus payment. The red solid line highlights the zero value. Figures 6a and 6b show the daily data between 12th January and 11th November 1919. Figures 6c and 6d show the semi-monthly series calculated using the daily data between 12th January and 11th November 1919. All figures are based on the 29 THBS households that used any credit purchases.
@@
Next, I look into consumers’ credit purchase behavior using the THBS dataset. Most of the THBS households ($29/33$) used any credit purchases during their sample period, supporting the evidence that this was a common practice among factory workers in Tsukishima. Figure 6a illustrates the average daily credit purchases and payments. Generally, there is a peak per semi-month in payments, which reflects the households’ behavior of making the payments for the credit purchase on or soon after their paydays. The purchases series shows a similar tendency but with lower spikes and larger variance. This indicates that the households purchased goods by a certain amount around their paydays and then used it when they needed unconservable food such as fish, fruit, and vegetables. Figure 6b illustrating the net purchases confirms this behavior: the households used the credit on a daily basis and redeemed it near paydays. Figure 6c shows that the payment slightly exceeds purchase, suggesting that many households could more or less repay their loans. Although the difference between payment and purchase is relatively obvious after August 1919, Figure 6d indicates that this difference is not statistically significant in all the semi-months. [^80] Strictly, payment can include accrued liabilities that should have been paid prior to the previous period. It is, therefore, reasonable that the average payment is greater than the average purchase in most semi-month cells. Thus, this is not a systematic trend but may be due to the smaller available number of households in the later sample period (Online Appendix Figure B.4).
I then analyze whether the volatility of earnings restricts credit purchases. Since the credit of a worker is unobservable, I use the volatility of earnings as a proxy of credit risk in the following empirical specification:
@@
where $I(\cdot)$ is an indicator function, $\overline{CP}$ is the (unit-)average credit purchases, $CV$ is the coefficient of variation for head’s earnings, $\bar{\mathbf{x}}$ is a vector of the (unit-)average family composition variables, and $v$ is a random error term. All variables are calculated using the semi-monthly series. [^81] For example, the coefficient of variation for household $i$ is defined as $CV_{i}=\{\sum_{t}^{T_{i}}(\tilde{y}_{i,t}-\bar{\tilde{y}}_{i,.})/(T_{i}-1)\}^{\frac{1}{2}}/\bar{\tilde{y}}_{i,.}$, where $T_{i}$ indicates the number of time-series observations, and $\bar{\tilde{y}}_{i,.}$ indicates the time-series average of the head’s income. In this cross-sectional model, I include a set of family composition variables to control for unobservable preferences because the household-fixed effect is unable to be modeled as an error component. If retailers allowed households to use credit purchases regardless of the variability of earnings, the estimate of $\varkappa_{1}$ becomes close to zero. In contrast, if retailers did not allow households with high-income volatility to trade on credit, the estimate should be positive and reasonably large.
@@
Notes: The dependent variable in Columns (1)-(3) is the indicator variable for the households who did not use any credit purchases during the sample period (non-CP users). The dependent variable in Columns (4)-(6) is the indicator variable for the non-CP users and households less than the first quartile in the sample average of credit purchases. Columns (1)-(2) and (4)-(5) use the linear probability model (LPM), whereas Columns (3) and (6) use the Probit model. “CV of Head’s Earnings” indicates the coefficient of variations in the head’s semi-monthly earnings. The sample mean of the CV is $5.76$ (Std. Dev. = $5.72$). All the regressions include the family composition variables: average family size, share of children aged 6–12, share of children aged 13–16, and share of adults aged 17+ (share of children aged 0–5 is used as a reference group). The average head’s age and average head’s earnings are added to the control variables in Columns 3 and 6. $F$-statistics $p$-values for the null of the zero-slope hypothesis are reported in Columns (1)-(2) and (4)-(5). $\chi^{2}$-statistics $p$-values for the null of the zero-slope hypothesis are listed in Columns (3) and (6). The standard errors for the LPM are based on the HC2 estimator proposed by Horn et al. (1975).
@@
Column (1) of Table 10 presents the result for equation 4. The estimated coefficient is close to zero and statistically insignificant. A potential concern herein may be the correlations between the volatility of earnings and the head’s age and earnings level. Column (2) shows the result from the extended model, including the head’s age and earnings in the controls, which shows the robustness of the baseline estimate in Column (1). Another potential issue is the functional form assumption in the linear probability model. Column (3) then lists an alternative non-linear model, including the full set of control variables used in Column (2). It confirms that the estimate is not statistically significantly different from zero. Next, I changed the threshold for the dependent indicator variable to consider a wider range of households using a relatively smaller amount of credit purchases, say less than the first quartile of the average credit purchases. Columns (4)-(5) summarize the results in the same column layouts as Columns (1)-(3), confirming that the estimates are still close to zero in all the specifications. Finally, the final row in this table shows that the null of the joint zero slope hypothesis is strongly rejected in all the regressions. Overall, there is no apparent trend in the characteristics of households to use credit purchases. This suggests that retailers may not have put weight on the volatility of earnings when trading with factory workers on credit. This is consistent with the retailers’ motivation to secure their local customers discussed earlier, and with the fact that factory workers commonly used credit purchases. [^82] Despite this, low-income households might have had more difficulty using credit purchases. A famous essay drawn by a primary school student in the 1930s depicts an example of a rice retailer who refused to buy rice under credit purchase (Toyoda 1969, pp. 346–355).
@@
The results so far indicate that retailers have been generous in the earnings volatility risk. This, in turn, implies that retailers may factor risk in the form of interest into their loans or retail prices. Unfortunately, there are no systematic statistics on the interest in credit purchases. Thus, I calculate the risk premiums backward by differencing payments and purchases. This backward calculation technique requires information both on the price and type of retailers. However, budget books usually document price-by-item information but not price-by-retailer information. This is problematic because the same items were often sold in different types of retailers: for example, eggs were sold in greengrocers (yaoya) in many cases but sometimes sold in variety stores (aramonoya). Fortunately, I found that a specific month in a budget book has documented the price-by-retailer information. [^83] The OISR, Archives of the Tsukishima Survey (THBS #4, February 1919). According to this budget book, the reimbursements were made semi-monthly, as seen in Figure 6b. They bought various goods in greengrocers, fish stores, and variety stores between 1st and 14th February 1919 and paid exactly the same amounts on the 15th. [^84] While the retailers used credit purchases from wholesalers (footnote 78), the repayment period was often set around one month (Tokyo City Office 1932, pp. 157-159). Therefore, this makes sense for consumers to make repayments at semi-monthly or monthly intervals. In the latter half of February, the household paid the same amounts to greengrocers and variety stores on 28th, but less amount to the fish stores. Since the budget book for March does not have information on the type of retailers, it is difficult to know exactly how the reimbursement shortfall was handled. However, this case suggests that the risk premiums may have been included in the retail prices but not in the form of interest.
@@
Finally, I assess the loss for the retailers in credit purchases. According to the statistics on the losses on credit sales documented in the 1930 Survey Report, almost all of the retailers had suffered losses in credit purchases. For example, 99% of rice and 97% of fuel retailers more or less lost their sales. This means that credit purchase was inherently a loss-making sales method for the retailers. Since systematic statistics are unavailable to clarify the percentage of total profits corresponding to these losses. Thus, it is impossible to discuss how severe these losses are for the retailers. However, an important fact is that these losses were usually capped by 5% of total sales. For example, the available statistics suggest that 87% of rice and 65% of fuel retailers suffered losses amounting to 5% of sales. This supports the evidence that while the retailers could not avoid losses in credit sales, they might have set prices that anticipated predicted losses to some extent. This seems to be a plausible sales strategy for retailers to deal with adverse selection issues under imperfect consumer information. Online Appendix C.4 summarizes the statistics on the losses in credit purchases available from the 1930 Survey Report in detail.
@@
To summarize, the available documents show that retailers competed with other retailers in the same sales area. Under this condition, they provided regular consumers with the opportunity to trade on credit without interest for at least one month. The 1930 Survey Report indeed stated that “intense competition among competitors has led to a high frequency of credit purchases” (Tokyo City Office 1932b, p. 415). This accessibility to credit purchases could help workers manage consumption after unexpected individual setbacks in the short term. [^85] Another potential explanation for the popularity of credit purchases might be thefts. However, the number of thefts $100$ people in 1919 is $0.5$ in Tsukishima, which is close to or slightly lower than the city average of $1.3$ (National Police Agency 1920b, pp 53–54; 244–245). In fact, Tsukishima Police Department does not mention that Tsukshima was an area with a particularly high rate of thefts (Tsukishima Police Department 1976, pp. 244–245). This suggests that this crime channel is not relevant to explain the use of credit purchases.

### v10

@@ 4.3 Event-Study Analysis: Response to Adverse Health Shocks
Column 1 shows the results of equation 2, in which the head’s earning is replaced with the head’s illness (i.e., the number of days with illness in the previous semi-month cell). The family size control, household fixed effects, and time-fixed effects are included in all the regressions. The estimated coefficients on the head’s illness are listed in the sub-column named “Coef.”.
Column 2 shows the results of equation 2, in which the head’s earning is treated as the endogenous variable. The head’s illness is an instrumental variable (i.e., excluded exogenous regressor) in the first-stage reduced-form equation. The first-stage result is identical to the result listed in Column (2) of Table 6 with the first-stage $F$-statistics of $13.6$. The estimated coefficients on the head’s earnings are listed in the sub-column named “Coef.”.
@@
Panel B: “Other member’s earnings” refers to the total income earned by all family members except for the head of the household. “Sales of miscellaneous assets” includes the sale of daily miscellaneous goods such as newspapers and empty bottles. All dependent variables are in yen.
@@ 5 Credit System in Local Economy
The foregoing results enlighten the roles of credit institutions among skilled workers’ households during industrialization. As the credit for purchases was provided by local retailers, this suggests the importance of the relationships between the workers and retailers. To the best of my knowledge, however, there is little historical research on credit purchases in prewar Japan. Hence, the details of this credit institution are not yet clear.
@@
Notes: This table summarizes the statistics on the sales in the $326$ retailers selling white rice, fish, fruit and vegetables, and firewood and charcoal in Tokyo surveyed in 1935. The number of stores reported the share of regular customers (column 2) is $90$, $56$, $61$, and $93$ in the white rice, fish, greengrocer, and fuel categories, respectively. The “share of sales for regular customers” is defined as the percentage of sales for regular customers in the annual sales. The definition of the social classes (“Upper”; “Middle”; “Other”) is not defined in the report. The share of sales method is based on the percentage of sales using each method in the annual sales. The number of competitors indicates the number of peers within a three-cho, approximately $327$ meters, radius. The department stores and retail markets are not included in this survey.
@@
The information on the credit purchases is personal, as in the case of state-contingent loans in the rural credit market (Udry 1994) and thus, not well documented. Therefore, systematic information on credit purchases is unavailable. Fortunately, however, I found that some official survey reports documented information on this credit institution. A survey report on retailers in 1930 (hereafter 1930 Survey Report) revealed that credit purchase was a popular payment method in transactions with consumers. [^71] This survey was conducted in February and March 1930 by Tokyo City and covered $3,892$ small and medium enterprises (Tokyo City Office 1932a, p. 1). Those included retailers as well as wholesale and other small businesses. The number of respondents varied depending on the question items. The number of respondents in the statistical table on the payment methods was $2,654$, of which $1,197$ were retailers (Tokyo City Office 1932b, p. 273). To the best of my knowledge, this is the most comprehensive survey of the retailers in Tokyo circa 1930. In fact, $99$%($155/157$) of grain retailers, $81$%($35/43$) of fish retailers, $78$%($64/82$) of greengrocers, and $94$%($45/48$) of fuel retailers used credit purchases for their selling method (Tokyo City Office 1932b, pp. 260; 263). A survey of retailers in 1935 published by the Tokyo City Chamber of Commerce and Industry provides more detailed information. Table 9 summarizes the information on the retailers’ sales measured in this survey. [^72] This survey covered $939$ retailers, and its report was edited on December 31, 1935. I digitized the statistics for all the retailers selling the representative items (rice, fish, fruit and vegetables, and fuels). The survey instrument included detailed questions regarding sales and purchases (Tokyo City Chamber of Commerce and Industry 1937, pp. legend, questionnaire). Online Appendix Table C.1 decomposes the statistics listed in Table 9 by scale of businesses. I confirm that the main findings are materially similar. First, many annual sales were from their regular customers, most of whom were from the non-upper classes. For example, $73$% of annual sales of rice retailers were from their regular customers, and more than $80$% of them were from non-upper classes, including factory workers. Second, credit purchase was a popular sales method, especially among rice and fuel retailers—more than $70$% of the sales were conducted with credit in both shops. [^73] These figures are systematically lower than those from the 1930 Survey Report. Given that both surveys show a systematic tendency toward credit purchases, this may be due to the differences in the samples, the definition of the retailers in each category, and the definition of the credit purchase itself. Although cash was also used among greengrocers, $34$% of the sales were settled using credit. This suggests that credit purchase was a popular payment method and that regular customers settled a large part of it. [^74] The 1930 Survey reveals that the retailers also commonly used credit purchases to buy their source goods from wholesalers (Tokyo City Office 1932a, pp. 154–157). This means that credit purchase was a familiar payment method for retailers as well as consumers. Although these statistics are from the mid-1930s, this confirms the possibility that the retailers were engaged in credit transactions with long-term relationships with their regular customers. [^75] A textbook for retailers published in 1924 explains the psychology behind this long-term relationship—credit purchase was provided for the convenience of the customer’s budget but also for the “vanity” (status) of being regular customers (Shimizu 1924, pp. 115–116). The final column of this table further provides suggestive evidence for this. The survey asked retailers whether the sales area was competitive or not, and if yes, they were asked to report the number of competitors within a three-cho (approximately $330$ meters) radius. As shown, rice and fuel stores have two to three more competitors than the retailers selling fish, fruit, and vegetables. Given the relatively high utilization of credit in rice and fuel retailers, they may have responded to competition by allowing their customers to trade on credit.
@@
Notes: Figure 6a, and 6c illustrate the time-series plots of the average daily and semi-monthly credit purchase and payment, respectively. The blue and red dashed lines indicate the average credit purchases and payments, respectively. Figure 6b and 6d illustrate the time-series plots of the average daily and semi-monthly credit purchase minus payment. The red solid line highlights the zero value. Figures 6a and 6b show the daily data between January 12 and November 11, 1919. Figures 6c and 6d show the semi-monthly series calculated using the daily data between January and November 11, 1919. All figures are based on the 29 THBS households that used any credit purchases.
@@
Next, I examined the credit purchases included in the THBS dataset. Most THBS households ($29/33$) used any credit purchases during their sample period, confirming that credit purchase was a common practice among factory workers. Figure 6a illustrates the average daily credit purchases and payments. Generally, there is a peak per semi-month in payments, which reflects the practice of making the payments for the credit purchase on or soon after their paydays. The purchase series shows a similar tendency but with lower spikes and larger variance. This indicates that the households first purchased goods for a certain amount on or soon after their paydays, and then used purchased perishable foods such as fish, fruit, and vegetables when they needed them. Figure 6b illustrating the net purchase confirms this behavior: the households used the credit on a daily basis and redeemed it near paydays. Figure 6c shows that the payment slightly exceeded purchase, suggesting that many households could approximately repay their loans. Although the difference between payment and purchase is relatively obvious after August 1919, Figure 6d indicates that this difference is not statistically significant in all the semi-months. [^76] Strictly, payment can include accrued liabilities that should have been paid prior to the previous period. It is, therefore, reasonable that the average payment is greater than the average purchase in most semi-month cells. Thus, this is not a systematic trend but may be due to the smaller available number of households in the later sample period (Online Appendix Figure B.4).
I then analyzed whether the volatility of earnings restricted credit purchases. Since credit of worker was unobservable, I used the volatility of earnings as a proxy of credit risk in the following empirical specification:
@@
where $I(\cdot)$ is an indicator function, $\overline{CP}$ is the (unit-)average credit purchases, $CV$ is the coefficient of variation for head’s earnings, $\bar{\mathbf{x}}$ is a vector of the (unit-)average family composition variables, and $v$ is a random error term. All variables are calculated using the semi-monthly series. [^77] For example, the coefficient of variation for household $i$ is defined as $CV_{i}=\{\sum_{t}^{T_{i}}(\tilde{y}_{i,t}-\bar{\tilde{y}}_{i,.})/(T_{i}-1)\}^{\frac{1}{2}}/\bar{\tilde{y}}_{i,.}$, where $T_{i}$ indicates the number of time-series observations, and $\bar{\tilde{y}}_{i,.}$ indicates the time-series average of the head’s income. In this cross-sectional model, I include a set of family composition variables to control for unobservable preferences because the household-fixed effect cannot be modeled as an error component. If retailers allowed households to use credit purchases regardless of the variability of earnings, the estimate of $\varkappa_{1}$ becomes close to zero. By contrast, if retailers did not allow households with high-income volatility to trade on credit, the estimate should be positive and reasonably large.
@@
Notes: The dependent variable in Columns (1)–(3) is the indicator variable for the households who did not use any credit purchases during the sample period (non-CP users). The dependent variable in Columns (4)–(6) is the indicator variable for the non-CP users and households less than the first quartile in the sample average of credit purchases. Columns (1)–(2) and (4)–(5) use the linear probability model (LPM), whereas Columns (3) and (6) use the Probit model. “CV of Head’s Earnings” indicates the coefficient of variations in the head’s semi-monthly earnings. The sample mean of the CV is $5.76$ (Std. Dev. = $5.72$). All the regressions include the family composition variables: average family size, share of children aged 6–12, share of children aged 13–16, and share of adults aged 17+ (share of children aged 0–5 is used as a reference group). The average head’s age and average head’s earnings are added to the control variables in Columns 3 and 6. $F$-statistics $p$-values for the null of the zero-slope hypothesis are reported in Columns (1)–(2) and (4)–(5). $\chi^{2}$-statistics $p$-values for the null of the zero-slope hypothesis are listed in Columns (3) and (6). The standard errors for the LPM are based on the HC2 estimator proposed by Horn et al. (1975).
@@
Column (1) of Table 10 presents the result for Equation 4. The estimated coefficient is close to zero and statistically insignificant. A potential concern may be the correlations between the volatility of earnings and the head’s age and earnings level. Column (2) shows the result from the extended model, including the head’s age and earnings in the controls, which shows the robustness of the baseline estimate in Column (1). Another potential issue is the functional form assumption in the linear probability model. Column (3) then lists an alternative non-linear model, including the full set of control variables used in Column (2). It confirms that the estimate is not statistically significantly different from zero. Next, I changed the threshold for the dependent indicator variable to consider a wider range of households including the households with less than the first quartile of the average credit purchases. Columns (4)–(6) summarize the results in the same column layouts as Columns (1)–(3), confirming that the estimates are still close to zero in all the specifications. Finally, the last row in this table shows that the null of the joint zero slope hypothesis is strongly rejected in all the regressions. Overall, there is no apparent trend in the characteristics of households that use credit for their purchases. This suggests that retailers may not have considered the volatility of earnings when providing credit to factory workers. This is consistent with the retailers’ motivation to secure their local customers discussed earlier, and with the fact that factory workers commonly used credit for purchases. [^78] Despite this, low-income households might have had more difficulty using credit for purchases. A famous essay drawn by a primary school student in the 1930s depicts an example of a rice retailer who refused to buy rice under credit purchase (Toyoda 1969, pp. 346–355).
@@
The results so far indicate that retailers were generous in considering the earnings volatility risk. This, in turn, implies that retailers may factor risk into retail prices or by charging interest on their loans. Unfortunately, there are no systematic statistics on the interest charged on credit purchases. It may be possible to reverse calculate the risk premiums by differencing payments and purchases. This reverse calculation technique requires information on both the price and type of retailers. However, budget books usually document price-by-item information but not price-by-retailer information. This is problematic because the same items were often sold in different types of retailers: for example, eggs were frequently sold in greengrocers (yaoya) but sometimes also in general stores (aramonoya). Fortunately, I found a specific month in a budget book that has documented the price-by-retailer information. [^79] The OISR, Archives of the Tsukishima Survey (THBS #4, February 1919). According to this budget book, the reimbursements were made semi-monthly, as seen in Figure 6b. They bought various goods in greengrocers, fish stores, and general stores between February 1 and February 14, 1919, and paid exactly the same amounts on the 15th. [^80] The repayment period in credit purchases among retailers and wholesalers (footnote 74) was often set at approximately one month (Tokyo City Office 1932, pp. 157-159). Therefore, customers likely made repayments at semi-monthly or monthly intervals. In the latter half of February, the household paid the same amounts to greengrocers and general stores on the 28th, but paid less to the fish stores. As the budget book for March does not have information on the type of retailers, it is difficult to determine exactly how the reimbursement shortfall was managed. However, this case suggests that the risk premiums may have been included in the retail prices but not in the form of interest.
@@
Finally, I assessed the losses suffered by retailers for credit purchases. According to the 1930 Survey Report, almost all the retailers suffered losses for credit purchases. For example, $99$% of rice and $97$% of fuel retailers lost their sales. This means that credit purchase was inherently a loss-making sales method for the retailers. As systematic statistics are unavailable to clarify the percentage of total profits corresponding to these losses, it is impossible to assess how severe these losses were for the retailers. However, an important fact is that these losses were usually capped at 5% of total sales. For example, the available statistics suggest that $87$% of rice and $65$% of fuel retailers suffered losses amounting to $5$% of sales. This supports the evidence that while the retailers could not avoid losses in credit sales, they might have set prices that anticipated predicted losses to some extent. This seems to be a plausible sales strategy for retailers to manage adverse selection issues under imperfect consumer information. Online Appendix C.4 summarizes the statistics on the losses in credit purchases available from the 1930 Survey Report in detail.
@@
To summarize, the available documents show that retailers competed with other retailers in the same sales area. Under this circumstance, they provided regular consumers with the opportunity to trade on credit without interest for at least one month. The 1930 Survey Report indeed stated that “intense competition among competitors has led to a high frequency of credit purchases” (Tokyo City Office 1932b, p. 415). This accessibility to credit purchases could help workers manage consumption after unexpected individual setbacks in the short term. [^81] Another potential explanation for the popularity of credit purchases might be theft. However, the number of thefts in $100$ people in 1919 was $0.5$ in Tsukishima, which is close to or even slightly lower than the city average of $1.3$ (National Police Agency 1920b, pp 53–54; 244–245). In fact, Tsukishima Police Department does not mention that Tsukshima was an area with a particularly high rate of theft (Tsukishima Police Department 1976, pp. 244–245). Therefore, the crime channel is not relevant to explain the use of credit purchases.